OpenStudy (anonymous):

give me at least 3 examples of each kind of number?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

give an example of a "kind of number"

OpenStudy (anonymous):

example of real number.. for elemtary algebra?

OpenStudy (anonymous):

integers? rationals? natural numbers?

OpenStudy (anonymous):

any number you say will be real

OpenStudy (anonymous):

\[2, 4, 0, \frac{1}{2}, \pi. \sqrt{5}\]

OpenStudy (anonymous):

3 examples of a real number..for elementary algebra....

OpenStudy (anonymous):

3 examples please

OpenStudy (anonymous):

above are 6 examples

OpenStudy (anonymous):

ok thank you.

OpenStudy (anonymous):

what is real numbers?